OpenStudy (anonymous):

Which function below has the smallest slope? f(x) f(x) = 5x - 4 g(x) graph going through (0, 1) and (2, 5) h(x) x h(x) 1 14 4 15 7 16 f(x) g(x) h(x) The functions all have the same slope

OpenStudy (anonymous):

\[f(x) = 5x - 4 \] has slope \(5\) as it looks like \(y=mx+b\)

OpenStudy (anonymous):

the slope through (0, 1) and (2, 5) is \[\frac{5-1}{2-0}=\frac{4}{2}=2\]

OpenStudy (anonymous):

the slope between \((1,14)\) and \((4,15)\) is \[\frac{15-14}{4-1}=\frac{1}{3}\]

OpenStudy (anonymous):

so \(f\) has slope \(5\), \(g\) has slope \(2\) and \(h\) has slope \(\frac{1}{3}\)

OpenStudy (anonymous):

so its h

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thank you so much!